OpenStudy (anonymous):

Plzzz help me....... I'll be ur fan.....anyone!!!!!!!!!!!!!!! Two ocean beaches are being affected by erosion. The table shows the width, in feet, of each beach at high tide measured where 1995 is represented by year 0. Year Western Beach Dunes Beach number width (in feet) width (in feet) 0 100 20 5 90 45 10 80 70 11 78 75 12 76 80 15 70 95 Describe the patterns shown by the erosion data measurements shown for each of the beaches in the table. a. Between which years will the beaches have approximately the same width? b. Assuming these rates remain constant, what can you do to get a better c. approximation of when the two beaches will have the same width?

OpenStudy (anonymous):

@Vocaloid @princeharryyy

OpenStudy (mertsj):

The western beach is getting wider. The dunes beach is getting narrower. Between years 11 and 12 the beaches will have about the same width. To get a more accurate answer, write the equations that express beach width as a function of the number of years for each beach. Solve the system of equations.

OpenStudy (anonymous):

so how do u do the last part.........?? @Mertsj

OpenStudy (anonymous):

Have you made the equations already ?

OpenStudy (anonymous):

no....

OpenStudy (anonymous):

You need to create the equations first to get a better approximation of when they will be the same

OpenStudy (anonymous):

Do you know how to create equations?

OpenStudy (anonymous):

ya...wait...

OpenStudy (anonymous):

western - y = -2x +100 dunes - y = 5x+20

OpenStudy (anonymous):

\[-2x +100 = 5x +20\] \[-2x+80=5x\] \[80=7x\] \[\frac{ 80 }{ 7 }=\frac{ 7x }{ 7 }\] \[11\frac{ 3 }{ 7 }\]

OpenStudy (anonymous):

Yes that is correct. In 11 3/7 months the two dunes will be equal

OpenStudy (anonymous):

ok thanks..........................